Some One please help to solve this
There are 12 points in a plane of which 5 are collinear. The number of triangle is
The answer Is 210
I don’t know how to do ???? please Explain


Answer by sudhanshu_kmr(1152)


To form a triangle we have to choose only 3 points in a plane but, these must not
collinear. So, firstly we will find total no. of ways to select 3 points from 12
points and then subtract the number of triangle that will form by using only 5
collinear points.

12C3 - 5C3 = 220- 10 = 210

if any doubt, u r welcome to contact :
http://www.algebra.com/tutors/aboutme.mpl?userid=sudhanshu_kmr